NatNah, a builder of acoustic accessories, has no debt and an equity cost of capital of 13%. Suppose NatNah decides to increase its leverage to maintain a market debt-to-value ratio of 0.5. Suppose its debt cost of capital is 8% and its corporate tax rate is 35%. If Natahs pre-tax WACC remains constant, what will be its (effective after-tax) WACC with the increase in leverage? The effective after-tax WACC will be %. (Round to two decimal places.)
Answer #1

For this, we'd use the MM Proposition II With tax. which is :-

Re = Ro + (R0 - RD) (1-t) D/E

Where Re the new cost of capital considering the effect of debt, which we have to calculate, and R0 is the Cost of capital without debt, RD is the cost of debt and D/E is debt to equity ratio.

Debt to value is debt/Total capital, that means debt is 50% of total capital hence the other 50% is equity, and the D/E Ratio is 1.

Ro = 13%

RD = 8%

(1-t) = (1-0.35) = 0.65

Putting everything in the formula above.

= 13% + (13-8)(0.65)(1)

= 16.25%

Hence the new effective after tax wacc is 16.25%.